OpenStudy (anonymous):

Integrate cos(2theta)d(theta) from 0 to pi/4

OpenStudy (anonymous):

\[\int\limits_{0}^{\Pi/4}\cos(2 \theta) d \theta\]

OpenStudy (anonymous):

\[ \int\cos(2\theta)d\theta =\frac{\sin(2\theta)}{2}\\ \int\limits_{0}^{\Pi/4}\cos(2 \theta) d \theta=\frac{\sin(2\cdot\frac{\pi}{4})}{2}-\frac{\sin(2\cdot 0)}{2} \]

OpenStudy (anonymous):

why is it divided by 2? :o

OpenStudy (anonymous):

Chain rule. If you derive \(\dfrac{\sin(2\theta)}{2}\) you'll get back \(\cos(2\theta)\). Without the division by 2 you would get back \(2\cos(2\theta)\).

OpenStudy (anonymous):

ooohhhhh okay ^.^ what about cos^2theta dtheta? is it cos^3theta/3?

OpenStudy (anonymous):

\[ \int\cos^2x\ dx=\int\frac{1+\cos2x}{2}dx=\frac{x}{2}+\frac{\sin2x}4 \] It isn't \(\cos^3x/3\) because if you derive that, you get \(\sin x\cdot\cos^2x\) because of the chain rule.

OpenStudy (anonymous):

is that an identity? cos^2x = 1+cos2x/2? :o

OpenStudy (anonymous):

Yep, it's an identity. You can derive it by combining the following two identities:\[ \sin^2x+\cos^2x=1\\\cos2x=\cos^2x-\sin^2x \]

OpenStudy (anonymous):

siiiick, thank you!!